a rectangle has a length of 17 inches more than three times its height if the perimeter is 498 what are the deminsions
Answer by rfer(16322)

2(3x+17)+2x=498
6x+34+2x=498
8x=498-34
8x=464
x=464/8
x=58
3x+17=191